# A Study of Fairness Concerns in AI-based Mobile App Reviews

### v2

@@ 3.1.2. Collection of potential fairness reviews
1. Search with an initial keywords set. While fairness is a confusing term and can have different meanings in different contexts (Ekstrand et al., 2022), our observation of exploring different literature reviews (e.g., (Xivuri and Twinomurinzi, 2021; Pessach and Shmueli, 2020; Mitchell et al., 2021; Mehrabi et al., 2021; Chen et al., 2022)) and grey literature (e.g., white papers issued by governments and institutions (Dawson et al., 2019; Anderson, 1992; MSR, [n. d.])) on fairness shows that ‘fair’, ‘discrimination’, and ‘bias’ are three commonly used keywords to refer to fairness. We performed a search using these three keywords on our set of 9,475,506 reviews, which resulted in a set of 30,019 ‘potential fairness reviews’.
@@
2. Using KeyBERT. ‘Fair’, ‘discrimination’, and ‘bias’ are not the only keywords that users use to discuss a fairness concern. Hence, we used KeyBERT (Grootendorst, 2020), a BERT-based keyword/phrase extraction technique, to extract more relevant fairness-related keywords and phrases, thereby increasing the representativeness and comprehensiveness of the final dataset. There are many embedding models that can be used in KeyBERT. We decided to use the all-mpnet-base-v2 embedding model as it is known to work well with English texts (Reimers, 2022). KeyBERT with the initial keywords (‘fair’, ‘discrimination’, and ‘bias’) as seed input was applied to the 30,019 potential fairness reviews to extract uni-gram, two-gram, and three-gram keywords. Given that the 30,019 reviews cover different topics, adding the seed keywords guided KeyBERT to look for keywords that are most similar and best describe fairness concerns in reviews.
KeyBERT suggested 210 keywords, including three initial keywords. In the next step, the second and third authors manually checked these keywords. They removed duplicates, the suggested 2-gram or 3-gram keywords (e.g., “fair policy”) that we had their uni-gram keywords (e.g., “fair”), and keywords that were rarely found in reviews (e.g., less than 0.001% of 9,475,506 reviews). This step resulted in reducing the number of keywords from 210 to 42 keywords. To make sure these 42 keywords are relevant and do not lead to unmanageable false positives, we next randomly selected 20 reviews per keyword matching, i.e., $42*20=840$ reviews. We removed duplicates in 840 reviews and reached 805 reviews. It was because some keywords appeared in more than one review. The second and third authors independently labelled these 805 reviews and classified them into 404 fairness reviews and 401 non-fairness reviews (see Fig. 2). They had 126 disagreements in this step, which were resolved using the negotiated agreement approach (Campbell et al., 2013). In this process of manual classification, we figured that 21 of these 42 keywords were irrelevant or redundant. This was determined by either (i) none of the matched reviews using these keywords was labelled as fairness review, or (ii) the matched reviews of another keyword subsumed all reviews matched by this keyword. For example, one of the irrelevant keywords was “suspend”. A review that matched this keyword was “My drive got suspended because of a DMCA (Digital Millennium Copyright Act) claim and I wasn’t active. I don’t have access to my personal data.” This review was labelled non-fairness review because it does not discuss anything about fairness. Next, we searched the remaining 21 keywords on 9,475,506 reviews and got 46,700 potential fairness reviews.
@@ 3.1.3. Manual labeling
The previous step resulted in 404 fairness reviews, 401 non-fairness reviews, and 46,700 potential fairness reviews. The 46,700 potential fairness reviews come from 108 apps. Note that these 108 apps possessed varied numbers of potential fairness reviews. Our aim was to build a reasonable and roughly balanced and diverse dataset of potential fairness reviews that reflect users’ opinions of different apps from various categories. To this end, we decided to randomly choose 45 potential fairness reviews from the apps that had at least 45 potential fairness reviews. This step resulted in 1,800 potential fairness reviews from 40 apps. This number of reviews (1,800 reviews) is statistically representative as it well exceeds the number needed for a statistical sample with a 99% confidence level and a 3% margin of error (Ahmad and Halim, 2017).
The labelling process of 1,800 potential reviews was conducted in several rounds and included five coders (i.e., five authors) with extensive experiences in human and social aspects of software engineering. This helped avoid fatigue, gradually reach a common understanding, and better manage disagreements. In each round, $\approx$300 reviews were independently labelled by two coders. In total, 227 disagreements were found between the coders. In case of any disagreement between two coders, a third coder was asked to read the review and label it. Hence, the final label was determined based on majority opinion. Our labelling process led to 728 fairness and 1,072 non-fairness reviews, which with 404 fairness and 401 non-fairness reviews in Section 3.1.2 constitute a dataset (ground-truth) of 1,132 fairness and 1,473 non-fairness reviews. Those potential fairness reviews labelled as non-fairness mainly reported a bug or did not discuss any concrete fairness concerns (e.g., “My video is not saving in my Gallery and it’s stopping at 80% then saved in drafts only. This is not so fair” and “I don’t think it’s fair”).
4. Fairness Classifier
@@ 4.1. Approach
This section discusses the development of our binary classifier that accurately distinguishes fairness reviews from non-fairness reviews. For this purpose, we decided to use a wide range of supervised classifiers, including classical ML classifiers and state-of-the-art (SOTA) pre-trained natural language models (LMs). We opted to experiment with various classifiers because (1) it was unknown to us from the beginning which classifier could produce the best result and the feature-based ML techniques have been known to outperform SOTA LMs (Ezzini et al., 2022); (2) although the literature has shown the promising performance of ML and DL classifiers in text classification tasks, their performance could vary on different text-based datasets (Aggarwal and Zhai, 2012; Caruana and Niculescu-Mizil, 2006; Abdalkareem et al., 2020); and (3) classifiers use different strategies to handle overfitting and explainability, and have different execution times (Kotsiantis et al., 2006; Abdalkareem et al., 2020).
@@ 4.1.1. Dataset
The dataset constructed in Section 3.1 was used to train and test our learning classifiers. The dataset contains 1,132 fairness reviews and 1,473 non-fairness reviews. Following the NLP best practices applied to app reviews (Dąbrowski et al., 2022), we cleaned the dataset by removing emojis, numbers, punctuation, and one-letter characters.
4.1.2. Feature Extraction Techniques
We used three common text feature extraction techniques, Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE), and Word2Vec in ML classifiers. We briefly describe these feature extraction methods and the configuration we used to generate feature vectors from the reviews in our dataset. Note that the DL models use their own techniques to encode the reviews to the features.
TF-IDF calculates the significance of a word in a text corpus (Ramos et al., 2003). We configured TF-IDF to generate a feature union by capturing 4-, 5-, and 6-gram characters from the reviews. We also set the feature union to the top 50,000 features (Vogel and Meghana, 2020).
Word2Vec is a word embedding technique that captures the semantic and syntactic relationships between words in a large corpus of text (Mikolov et al., 2013a). We used the pre-trained 300-dimensional Google News Word2Vec embedding model containing 3M words and phrases, to produce a vector for each word in reviews of our dataset (Mikolov et al., 2013b).
USE works at the sentence level and utilizes pre-trained sentence embedding models to produce sentence vectors (Cer et al., 2018). We used the DAN encoder to encode a review into a 512-dimensional vector, as DAN encoders are better suited for large-scale tasks and are efficient regarding computational resources. As app reviews analysis, can be based on many reviews, we decided to go for DAN-based USE.
Combinations. Research has shown that using more than one feature extraction technique may help to yield better results in ML classifiers (Shahin et al., 2023). Hence, we also experimented with the various combinations of these three techniques to understand if their combinations enhance the performance of the ML classifiers.
4.1.3. Binary Classifiers
@@
In total, we used ten binary classifiers. Here, we describe six of them that performed better than the other four, due to space constraints. Previous studies (Aniche et al., 2020; Fu et al., 2016; Menzies and Shepperd, 2012) indicated that the parameters of these prediction models could be fine-tuned to improve their performance. Following the recommendations in (Bao et al., 2019; Tantithamthavorn et al., 2018), we checked several scenarios and values for the hyperparameters of each prediction model to find the best configuration for each prediction model. Specifically, we discuss how we implemented and fine-tuned them to achieve their best performance. The implementation and configuration of the rest of the classifiers can be found in our replication package (Nasab et al., 2024).
@@
Logistic Regression (LR) is a widely used classifier for binary classification tasks, which estimates the probability of a positive class given input features (Dreiseitl and Ohno-Machado, 2002). Considering the size of our dataset, we regularized the solver parameter to “liblinear” (Learn, 2007). Also, the penalty was set to L2, which could shrink the coefficients of the features and reduce overfitting.
@@
SVM (Support Vector Machine) uses mathematical principles to find the best way to classify data into different classes (Vapnik, 1999). In this study, we set the kernel parameter of this classifier to linear to compute the similarity of two input vectors.
@@
RoBERTa (Robustly Optimized BERT Pretraining Approach) is a modified version of BERT that enhances performance on various NLP tasks (Liu et al., 2019). As an input to RoBERTa, the reviews in our dataset were transformed to a sequence of vectors using the Hugging Face RoBERTa tokenizer that uses the byte-level Byte-Pair Encoding (BPE) algorithm and “roberta-base”, a pre-trained model with 125 million parameters (face, 2020). In both BERT and RoBERTa, we used linear transformation to map the hidden states to the output vocabulary size. Then a combination of the sigmoid layer and binary cross-entropy loss was used as the last layer in both models that predicts whether a review is a fairness review or not. We also used the Adam algorithm (Kingma and Ba, 2014) with a learning rate of 2e-5 to optimize the models and trained their networks for 10 epochs with a batch size of 15.
GPT-3 (Generative Pre-training Transformer) developed by OpenAI [^1] https://openai.com is a large language model trained on a huge amount of text data that can produce human-like texts for various tasks (Brown et al., 2020). Its performance can be further improved with human feedback (Brown et al., 2020). GPT-3 uses BPE (Byte-Pair Encoding) as a tokenizer to map a given text into a sequence of integers that represent the text. This process relies on “text-davinci-003”, a pre-trained model with 175 billion parameters built on top of the InstructGPT model (Brown et al., 2020).
@@ 4.1.4. Evaluation Metrics
Similar to other studies (Yang et al., 2022; Abdalkareem et al., 2020; Viviani et al., 2019), we used precision, recall, F1-score, accuracy, and AUC to show the performance of the classifiers. These five common metrics are calculated based on False Positive (FP), True Positive (TP), True Negative (TN), and False Negative (FN). In this paper, precision is the proportion of correctly predicted fairness reviews out of all reviews predicated as fairness reviews (i.e., $Precision=\frac{TP}{TP+FP}$). Recall is the proportion of correctly predicted fairness reviews out of the total number of fairness reviews (i.e., $Recall=\frac{TP}{TP+FN}$). F1-score combines precision and recall into one single score (i.e., $F1-score=\frac{2*Precision*Recall}{Precision+Recall}$). Accuracy is the ratio of precisely predicted fairness reviews and non-fairness reviews to all reviews (i.e., $Accuracy=\frac{TP+TN}{TP+FP+TN+FN}$). AUC measures the likelihood of a classifier ranking a randomly selected fairness review (i.e., TP) higher than a randomly selected non-fairness review (i.e., FP). The AUC provides an aggregate measure of performance across all possible classification thresholds. An AUC=1.0 signifies a perfect classifier (i.e., able to perfectly discriminate between the two classes), while an AUC=0.5 suggests that the classifier is no better than random guessing; a higher AUC value closest to 1.0 is desired.
We used the 10-fold cross-validation method to evaluate the classifiers except for GPT-3 (Efron, 1983). The 10-fold cross-validation method splits our dataset (ground-truth) into ten parts (i.e., fold), which are used for training and testing the classifiers ten times. In each iteration, nine parts (90% of the dataset) are used as training data and one part for evaluating the classifiers. We calculated the performance of GPT-3 by comparing the labels (fairness or not-fairness) generated by GPT-3 and the labels of the reviews in the dataset (ground-truth). We used the GPT-3 API and simply asked GPT-3 to determine if reviews in the dataset are about fairness concerns or not (i.e., by prompting “Tell me if the following text is related to a fairness concern or not. Just say Yes or No”).
Table 1. Results of the binary classifiers (in %). The best result in each metric is greyed and bold. The model chosen as the best-performing model is bold and yellowed. P: Precision, R: Recall, F1: F1-score, ACC: Accuracy, AUC: Area under the ROC Curve, XGB: XGBoost, W2V: Word2Vec.

### v3

@@ 3.1.2. Collection of potential fairness reviews
1. Search with an initial keywords set. While fairness is a confusing term and can have different meanings in different contexts (Ekstrand et al., 2022), our observation of exploring different literature reviews (e.g., (Xivuri and Twinomurinzi, 2021; Pessach and Shmueli, 2022; Mitchell et al., 2021; Mehrabi et al., 2021; Chen et al., 2022)) and grey literature (e.g., white papers issued by governments and institutions (Dawson et al., 2019; Anderson, 1992; MSR, [n. d.])) on fairness shows that ‘fair’, ‘discrimination’, and ‘bias’ are three commonly used keywords to refer to fairness. We performed a search using these three keywords on our set of 9,475,506 reviews, which resulted in a set of 30,019 ‘potential fairness reviews’.
@@
2. Using KeyBERT. ‘Fair’, ‘discrimination’, and ‘bias’ are not the only keywords that users use to discuss a fairness concern. Hence, we used KeyBERT (Grootendorst, 2020), a BERT-based keyword/phrase extraction technique, to extract more relevant fairness-related keywords and phrases, thereby increasing the representativeness and comprehensiveness of the final dataset. There are many embedding models that can be used in KeyBERT. We decided to use the all-mpnet-base-v2 embedding model as it is known to work well with English texts (Reimers, 2022). We conducted several pilot experiments to understand if KeyBERT should be applied to all 9M reviews or the 30,019 potential fairness reviews. Our experiments revealed that applying KeyBERT to all 9M reviews led to many irrelevant suggested keywords and their corresponding reviews. Hence, KeyBERT with the initial keywords (‘fair’, ‘discrimination’, and ‘bias’) as seed input was applied to the 30,019 potential fairness reviews to extract uni-gram, two-gram, and three-gram keywords. This approach produced less noise while identifying the most relevant keywords and fairness reviews. Given that the 30,019 reviews cover different topics, adding the seed keywords guided KeyBERT to look for keywords that are most similar and best describe fairness concerns in reviews.
KeyBERT suggested 210 keywords, including three initial keywords. In the next step, the second and third authors manually checked these keywords. They removed duplicates, the suggested 2-gram or 3-gram keywords (e.g., “fair policy”) that we had their uni-gram keywords (e.g., “fair”), and keywords that were rarely found in reviews (e.g., less than 0.001% of 9,475,506 reviews). This step resulted in reducing the number of keywords from 210 to 42 keywords. To make sure these 42 keywords are relevant and do not lead to unmanageable false positives, we next randomly selected 20 reviews per keyword matching, i.e., $42*20=840$ reviews. We removed duplicates in 840 reviews and reached 805 reviews. It was because some keywords appeared in more than one review. The second and third authors independently labeled these 805 reviews. Given the absence of a universally accepted definition of fairness, the second and third authors were asked to review various definitions of fairness from different domains. These definitions were collected from both peer-reviewed literature (Bird et al., 2020; Mehrabi et al., 2021; Pessach and Shmueli, 2022; d’Aloisio et al., 2024; Liu et al., 2024) and gray literature (Dawson et al., 2019; Anderson, 1992; MSR, [n. d.]; IEEE, 2021). For example, Pessach and Shmueli (Pessach and Shmueli, 2022) argued that the legal domain defines discrimination in two ways: direct discrimination, which occurs when an individual is intentionally treated differently based on his/her membership in a protected class, and indirect discrimination, which occurs when members of a protected class are negatively affected more than others even if by a seemingly neutral policy. This ‘definition review’ process was intended to guide their labeling decisions on whether a particular review should be classified as a fairness review. The inter-rater agreement between the second and third authors was measured using the Cohen’s kappa coefficient (McHugh, 2012), which yielded a value of 0.69. This indicates a substantial level of agreement. The two authors had 126 disagreements in this step, which were resolved using the negotiated agreement approach (Campbell et al., 2013). The main reason behind the disagreements between the two authors could be attributed to the fact that they have different understandings of fairness and the definitions of fairness provided to them. At the end of this step, the 805 reviews were classified into 404 fairness reviews and 401 non-fairness reviews (see Fig. 2).
In this process of manual classification, we figured that 21 of these 42 keywords were irrelevant or redundant. This was determined by either (i) none of the matched reviews using these keywords was labeled as fairness review, or (ii) the matched reviews of another keyword subsumed all reviews matched by this keyword. For example, one of the irrelevant keywords was “suspend”. A review that matched this keyword was “My drive got suspended because of a DMCA (Digital Millennium Copyright Act) claim and I wasn’t active. I don’t have access to my personal data.” This review was labeled non-fairness review because it does not discuss anything about fairness. Next, we searched the remaining 21 keywords on 9,475,506 reviews and got 46,700 potential fairness reviews. Inspired by the collected definitions of fairness and our observations in this step, we defined a fairness review as follows: a review in which users express that they are treated differently, whether intentionally or unintentionally, due to their inherent, acquired, or context-specific attributes (e.g., age, language, location, gender) by apps or the decisions and policies of the app owners. Additionally, a fairness review can reflect users’ perceptions that the outcomes of apps, or the decisions and policies of the app owners, are inconsistent, non-transparent, and unreasonable. In such fairness reviews, users do not often explicitly compare themselves with others based on inherent, acquired, or context-specific attributes. Instead, they typically complain about outcomes, decisions, and policies that they perceive as inconsistent, non-transparent, and unreasonable (e.g., “I used to love this app, but since last month, for no reason, my account has been getting restricted again and again every 24 hours. I am blocked. [the app name] should do something about it; that’s unfair for users like me who have never used any third-party app, bot, or spam and still suffer the restrictions and shadow ban.”).
@@ 3.1.3. Manual labeling
The previous step resulted in 404 fairness reviews, 401 non-fairness reviews, and 46,700 potential fairness reviews. The 46,700 potential fairness reviews come from 108 apps. Note that these 108 apps possessed varied numbers of potential fairness reviews. Our aim was to build a reasonable and roughly balanced and diverse dataset of potential fairness reviews that reflect users’ opinions of different apps from various categories. To this end, we decided to randomly choose 45 potential fairness reviews from the apps that had at least 45 potential fairness reviews. This step resulted in 1,800 potential fairness reviews from 40 apps. This number of reviews (1,800 reviews) is statistically representative as it well exceeds the number needed for a statistical sample with a 99% confidence level and a 3% margin of error (Israel et al., 1992; Ahmad and Halim, 2017).
The labeling process of 1,800 potential reviews was conducted in three rounds and included five coders with extensive experiences in the human and social aspects of software engineering. This helped avoid fatigue, gradually reach a common understanding, and better manage disagreements. In each round, $\approx$300 reviews were independently labeled by two coders. The five coders included the two coders who were involved in the previous step and three new coders (i.e., the second, fourth, and fifth authors). While the definition of fairness reviews was given to all coders, including new coders, to guide labeling their allocated reviews, they also accessed the various fairness definitions collected in the previous step. The inter-rater agreement between the coders was quantified using the Cohen’s kappa coefficient (McHugh, 2012), resulting in a coefficient of 0.75, which shows a substantial agreement. This level of agreement indicates high reliability in the coding process. In total, 227 disagreements were found between the coders. In case of any disagreement between two coders, a third coder was asked to read the review and label it. Hence, the final label was determined based on majority opinion. Our labeling process led to 728 fairness and 1,072 non-fairness reviews, which with 404 fairness and 401 non-fairness reviews in Section 3.1.2 constitute a dataset (ground-truth) of 1,132 fairness and 1,473 non-fairness reviews. Those potential fairness reviews labeled as non-fairness mainly reported a bug or did not discuss any concrete fairness concerns (e.g., “My video is not saving in my Gallery and it’s stopping at 80% then saved in drafts only. This is not so fair” and “I don’t think it’s fair”).
4. Fairness Reviews Identification (RQ1)
@@ 4.1. Approach
This section discusses the development of our model that accurately distinguishes fairness reviews from non-fairness reviews. For this purpose, we decided to use a wide range of classification models, including classical ML models, state-of-the-art (SOTA) pre-trained language models (LMs), and large language models (LLMs). We opted to experiment with various models because (1) it was unknown to us from the beginning which model could produce the best result and the feature-based ML techniques have been known to outperform SOTA pre-trained LMs (Ezzini et al., 2022); (2) although the literature has shown the promising performance of ML and DL models in text classification tasks, their performance could vary on different text-based datasets (Aggarwal and Zhai, 2012; Caruana and Niculescu-Mizil, 2006; Abdalkareem et al., 2020); and (3) classifiers use different strategies to handle overfitting and explainability, and have different execution times (Kotsiantis et al., 2006; Abdalkareem et al., 2020).
@@ 4.1.1. Dataset
The dataset constructed in Section 3.1 was used to train and test our learning models. The dataset contains 1,132 fairness reviews and 1,473 non-fairness reviews. Following the NLP best practices applied to app reviews (Dąbrowski et al., 2022), we cleaned the dataset by removing emojis, numbers, punctuation, and one-letter characters.
4.1.2. Models
@@
In total, we used ten models. Here, we describe six of them that performed better than the other four, due to space constraints. Previous studies (Aniche et al., 2020; Fu et al., 2016; Menzies and Shepperd, 2012) indicated that the parameters of these prediction models could be fine-tuned to improve their performance. Following the recommendations in (Bao et al., 2019; Tantithamthavorn et al., 2018), we checked several scenarios and values for the hyperparameters of each prediction model to find the best configuration for each prediction model. Specifically, we discuss how we implemented and fine-tuned them to achieve their best performance. The implementation and configuration of the rest of the models can be found in our replication package (Nasab et al., 2024).
Traditional ML Models. In the traditional ML models, we used three common text feature extraction techniques: Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE), and Word2Vec. Initially, we briefly describe these feature extraction methods and the configuration we used to generate feature vectors from the reviews in our dataset. Then, we explain the fine-tuning of the traditional ML models.
TF-IDF calculates the significance of a word in a text corpus (Ramos et al., 2003). We configured TF-IDF to generate a feature union by capturing 4-, 5-, and 6-gram characters from the reviews. We also set the feature union to the top 50,000 features (Vogel and Meghana, 2020). Word2Vec is a word embedding technique that captures the semantic and syntactic relationships between words in a large corpus of text (Mikolov et al., 2013a). We used the pre-trained 300-dimensional Google News Word2Vec embedding model containing 3M words and phrases, to produce a vector for each word in reviews of our dataset (Mikolov et al., 2013b). USE works at the sentence level and utilizes pre-trained sentence embedding models to produce sentence vectors (Cer et al., 2018). We used the DAN encoder to encode a review into a 512-dimensional vector, as DAN encoders are better suited for large-scale tasks and are efficient regarding computational resources. As app reviews analysis, can be based on many reviews, we decided to go for DAN-based USE.
Research has shown that using more than one feature extraction technique may help to yield better results in ML models (Shahin et al., 2023). Hence, we also experimented with the various combinations of these three techniques to understand if their combinations enhance the performance of the ML models.
@@
LR (Logistic Regression) is a widely used model for binary classification tasks, which estimates the probability of a positive class given input features (Dreiseitl and Ohno-Machado, 2002). Considering the size of our dataset, we regularized the solver parameter to “liblinear” (Learn, 2007). Also, the penalty was set to L2, which could shrink the coefficients of the features and reduce overfitting.
@@
SVM (Support Vector Machine) uses mathematical principles to find the best way to classify data into different classes (Vapnik, 1999). In this study, we set the kernel parameter of this model to linear to compute the similarity of two input vectors.
@@
Pre-trained Language Models. In this study, we used two widely used pre-trained models to detect fairness reviews automatically.
@@
RoBERTa (Robustly Optimized BERT Pretraining Approach) is a modified version of BERT that enhances performance on various NLP tasks (Liu et al., 2019). As an input to RoBERTa, the reviews in our dataset were transformed to a sequence of vectors using the Hugging Face RoBERTa tokenizer that uses the byte-level Byte-Pair Encoding (BPE) algorithm and “roberta-base”, a pre-trained model with 125 million parameters (face, 2020). We used linear transformation in both BERT and RoBERTa to map the hidden states to the output vocabulary size. Then, a combination of the sigmoid layer and binary cross-entropy loss was used as the last layer in both models to categorize a review into a fairness review or a non-fairness review. We also used the Adam algorithm (Kingma and Ba, 2014) with a learning rate of 2e-5 to optimize the models and trained their networks for 10 epochs with a batch size of 15.
Large Language Models (LLMs). Finally, we used GPT-3 (Generative Pre-training Transformer) developed by OpenAI [^1] https://openai.com to identify fairness reviews. GPT-3 is a large language model trained on a huge amount of text data that can produce human-like texts for various tasks (Brown et al., 2020). Its performance can be further improved with human feedback (Brown et al., 2020). GPT-3 uses BPE (Byte-Pair Encoding) as a tokenizer to map a given text into a sequence of integers that represent the text. This process relies on “text-davinci-003”, a pre-trained model with 175 billion parameters built on top of the InstructGPT model (Brown et al., 2020).
Table 2. The metrics used to evacuate the models. TP: True Positive, FP:False Positive, TN: True Negative, FN: False Negative, TPR: True Positive Rate, FPR: False Positive Rate
Metrics Calculations Descriptions
Precision $\frac{TP}{TP+FP}$ It is the proportion of correctly predicted fairness reviews out of all reviews predicated as fairness reviews.
Recall $\frac{TP}{TP+FN}$ It is the proportion of correctly predicted fairness reviews out of the total number of fairness reviews.
F1-score $\frac{2*Precision*Recall}{Precision+Recall}$ It combines precision and recall into one single score.
Accuracy $\frac{TP+TN}{TP+FP+TN+FN}$ It is the ratio of precisely predicted fairness reviews and non-fairness reviews to all reviews.
AUC $\int_{0}^{1}TPR(t_{i}),dFPR(t_{i})$ It measures the likelihood of a model ranking a randomly selected fairness review (i.e., True Positive Rate) higher than a randomly selected non-fairness review (i.e., False Positive Rate). The AUC provides an aggregate measure of performance across all possible model thresholds (i.e., $t_{i}$).
@@ 4.1.3. Evaluation Metrics
Similar to other studies (Yang et al., 2022; Abdalkareem et al., 2020; Viviani et al., 2019), we used precision, recall, F1-score, accuracy, and AUC to show the performance of the models. As shown in Table 2, these five common metrics are calculated based on False Positive (FP), True Positive (TP), True Negative (TN), and False Negative (FN). Note that AUC=1.0 signifies a perfect model (i.e., able to perfectly discriminate between the two classes), while an AUC=0.5 suggests that the model is no better than random guessing; a higher AUC value closest to 1.0 is desired (Berrar et al., 2019).
We used the 10-fold cross-validation method to evaluate the predictive capabilities of the traditional ML models and pre-trained language models (Efron, 1983). The 10-fold cross-validation method splits our dataset (ground-truth) into ten parts (i.e., fold), which are used for training and testing the models ten times. In each iteration, nine parts (90% of the dataset) are used as training data and one part for evaluating the models. We calculated the performance of GPT-3 by comparing the labels (fairness or not-fairness) generated by GPT-3 and the labels of the reviews in the ground-truth dataset. We used the GPT-3 API and simply asked GPT-3 to determine if reviews in the dataset are about fairness concerns or not (i.e., by prompting “Tell me if the following text is related to a fairness concern or not. Just say Yes or No”).
Table 3. Results of the models (in %). The best result in each metric is greyed and bold. The model chosen as the best-performing model is bold and yellowed. P: Precision, R: Recall, F1: F1-score, ACC: Accuracy, AUC: Area under the ROC Curve, XGB: XGBoost, W2V: Word2Vec.